OpenStudy (anonymous):

Find the exact value of each function for the given angle for f(θ) = sin θ and g(θ) = cos θ. Do not use a calculator. θ = 180° find (fg)(θ)

OpenStudy (anonymous):

\[ (fg)(\theta) = (\sin \theta)(\cos\theta) = (\sin 180^\circ) (\cos 180^\circ) \]

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!